# MAS-Bench: A Unified Benchmark for Shortcut-Augmented Hybrid Mobile GUI Agents

### v1

@@ Abstract
To enhance the efficiency of GUI agents on various platforms like smartphones and computers, a hybrid paradigm that combines flexible GUI operations with efficient shortcuts (e.g., API, deep links) is emerging as a promising direction. However, a framework for systematically benchmarking these hybrid agents is still underexplored. To take the first step in bridging this gap, we introduce MAS-Bench, a benchmark that pioneers the evaluation of GUI-shortcut hybrid agents with a specific focus on the mobile domain. Beyond merely using predefined shortcuts, MAS-Bench assesses an agent’s capability to autonomously generate shortcuts by discovering and creating reusable, low-cost workflows. It features 139 complex tasks across 11 real-world applications, a knowledge base of 88 predefined shortcuts (APIs, deep-links, RPA scripts), and 7 evaluation metrics. The tasks are designed to be solvable via GUI-only operations, but can be significantly accelerated by intelligently embedding shortcuts. Experiments show that hybrid agents achieve significantly higher success rates and efficiency than their GUI-only counterparts. This result also demonstrates the effectiveness of our method for evaluating an agent’s shortcut generation capabilities. MAS-Bench fills a critical evaluation gap, providing a foundational platform for future advancements in creating more efficient and robust intelligent agents. The project page is available at https://pengxiang-zhao.github.io/MAS-Bench.
@@ 1 Introduction
The rise of Large Language Models (LLMs) (et al. 2024; Zhao et al. 2023; Minaee et al. 2024; Chang et al. 2024; Xu et al. 2024) is driving the development of Graphical User Interface (GUI) Agents, enabling them to operate diverse digital platforms such as computers, web browsers, and smartphones (Liu et al. 2025b; Tang et al. 2025; Liu et al. 2025a). Early research on mobile agents primarily focused on replicating human-like flexibility through GUI-only interaction (Lee et al. 2025; Gou et al. 2024; Cheng et al. 2024; You et al. 2024; Zhang et al. 2024; Lu et al. 2025). This approach grants agents the generality to operate on any application, but it often overlooks the significant efficiency advantages offered by more direct methods.
As GUI agents are increasingly applied to complex tasks, enhancing their efficiency has become a central research focus. In this context, hybrid paradigms have emerged as a promising solution, demonstrating effectiveness across a wide range of platforms (Wang et al. 2025b; Zheng et al. 2025; Zhang et al. 2025a, b; Jiang et al. 2025; Lee et al. 2023). This approach combines the speed and reliability of “shortcuts” such as Application Programming Interface (API) calls, deep links, and Robotic Process Automation (RPA) scripts (Bridle and McCreath 2006; Guerreiro, Gamboa, and Jorge 2008; Kennedy and Everett 2011; Tripathi 2018; Ling, Gao, and Wang 2020; Agostinelli et al. 2022), with the flexibility of GUI operations. As Fig. 1 illustrates, a hybrid agent can bypass tedious, multi-step GUI operations by invoking a single, efficient shortcut, drastically reducing operational complexity and time.
@@
However, despite these advancements, a framework for systematically evaluating and benchmarking hybrid agents is still lacking (Le et al. 2020; Roffarello, Purohit, and Purohit 2024; Zhang et al. 2025a, b). This gap leaves the full potential of GUI-Shortcut agents far from being thoroughly explored and assessed. While this is a general challenge, the mobile domain represents a critical and demanding environment for progress.
Therefore, we present MAS-Bench, the first benchmark explicitly designed to evaluate GUI-shortcut hybrid mobile agents. The tasks within MAS-Bench are solvable entirely through GUI interactions, but integrating shortcuts can improve efficiency. MAS-Bench evaluates agents’ ability to make optimal decisions between different operational modes to enhance performance. Our experiments on MAS-Bench show that hybrid agents achieve up to a 64.1% success rate, a significant improvement over the 44.6% achieved by their GUI-only counterparts, and execute tasks with over 40% greater efficiency.
Furthermore, we introduce a novel framework to evaluate agents’ capacity to generate new shortcuts from interaction. This framework integrates agent-generated shortcuts into a standard baseline agent and measures subsequent task performance. Our findings reveal a performance gap: while our predefined shortcuts prove highly reliable (100% success rate), agent-generated shortcuts lag in robustness and efficiency. Dynamic shortcuts demonstrate significant potential. Despite their task completion rate of 38% (baseline 43%), they offer the highest efficiency, highlighting them as a promising direction for future research into robust shortcut generation. Our contributions are threefold:
  • We introduce MAS-Bench, the first benchmark for systematically evaluating GUI-shortcut hybrid mobile agents. It comprises 139 complex tasks spanning 11 real-world applications, supported by a knowledge base of 88 predefined shortcuts and 7 distinct evaluation metrics.
  • We establish baselines with several agents on MAS-Bench, demonstrating that those leveraging GUI-shortcut hybrid operations achieve significantly higher success rates and efficiency than their GUI-only counterparts.
  • We propose the first framework to evaluate an agent’s ability to generate shortcuts from interaction. Our experiments reveal a substantial gap between agent-generated shortcuts and predefined ones in efficiency and robustness, highlighting a key area for future research.
@@ 2 Related Work
Mobile task automation evolves from methods based on predefined scripts to more intelligent and adaptive agents driven by LLMs (Kirubakaran and Karthikeyani 2013; Amalfitano et al. 2014; Linares-Vásquez, Moran, and Poshyvanyk 2017; Kong et al. 2018; Zhao, Harrison, and Yu 2024). Traditional approaches, such as API calls, deep links, and Robotic Process Automation (RPA) scripts, offer direct execution paths but suffer from significant limitations, including invalidation from app updates, and a sensitivity to UI changes that hinders their ability to adapt (Bridle and McCreath 2006; Guerreiro, Gamboa, and Jorge 2008; Kennedy and Everett 2011; Le et al. 2020).
The rapid development of LLM-based GUI agents significantly overcomes traditional automation challenges by simulating human interaction through visual or multimodal inputs (Gou et al. 2024; Cheng et al. 2024; Baechler et al. 2024; Liu et al. 2025a). However, complete reliance on LLMs for fine-grained GUI interaction presents new problems: step-by-step action for routine tasks leads to inefficiency and increased costs; in complex operations, LLM hallucinations or misinterpretations can cause cumulative errors, affecting task success rates and reliability (Wen et al. 2024; Wang et al. 2024a). Zhang et al. (Zhang et al. 2025a) provide a detailed analysis of API and GUI agents; the former gain attention for their stability and programming integration capabilities, while the latter possesses strong adaptability. These considerations regarding the efficiency and robustness of GUI agents lead to an exploration of how to combine LLM with efficient shortcut methods, thereby promoting the rise of research in GUI-Shortcut agents.
@@
Researchers explore GUI-Shortcut methods to enhance agent capabilities and address the persistent challenges in task completion rate, costs, and operational efficiency for LLM-based GUI agents. For instance, UFO2 (Zhang et al. 2025b), a desktop AgentOS, demonstrates robust task execution through its multi-agent architecture and a unified GUI-API action layer that deeply integrates with operating system and application features. Inspired by such work, AppAgentX (Jiang et al. 2025) improves efficiency by evolving high-level actions into shortcuts based on task execution history. MobileAgent-E (Wang et al. 2025a) uses its self-evolving module, learns and stores general-purpose Tips and reusable Shortcuts from past experiences to continuously improve performance and efficiency. While these efforts focused on the mobile domain are promising, they also highlight the urgent need for a systematic method to evaluate the effectiveness of these emerging hybrid operations on mobile devices.
@@
AndroidArena 221
AndroidWorld 116
LlamaTouch 495
AndroidLab 138
SPA-Bench 340
@@
Mobile GUI Agent Benchmark.
As agents evolve from GUI-based interaction to hybrid GUI-Shortcut operations, they place new demands on benchmarking. Among existing benchmarks, MobileAgentBench (Wang et al. 2024c) focuses on the foundational GUI navigation and task completion capabilities of LLM-powered mobile agents. Mobile-Eval-E (Wang et al. 2025a) includes complex, real-world mobile tasks that require long-horizon, cross-app interactions. While these benchmarks advance research in various dimensions, there remains a need for a benchmark that can comprehensively evaluate the ability of agents to intelligently discover, decide upon, and execute diverse shortcuts within GUI interactions and also measure the overall task effectiveness of this GUI-Shortcut hybrid operation (Zhang et al. 2025b).
@@ 3.1 Online Evaluation Environment
MAS-Bench is built on a dynamic Android platform to evaluate mobile agent performance on complex tasks comprehensively. Unlike static benchmarks, which are often limited to fixed datasets, our platform enables the real-time assessment of agents’ capacity to select and utilize shortcuts intelligently. This dynamic approach moves beyond simply verifying task outcomes, focusing on the agent’s decision-making process in live scenarios. To ensure real-world relevance, the benchmark incorporates a suite of widely used Android applications covering diverse daily-life scenarios. All evaluations are conducted on a standardized emulator to ensure fairness, eliminating confounding factors such as device variations and environmental inconsistencies.
@@ 3.2 GUI-Shortcut Hybrid Action Space
To complete tasks, agents in MAS-Bench operate within a hybrid action space that combines conventional GUI interactions with programmatic shortcuts. This space comprises two primary action types:
@@
  • API: Direct programmatic control that executes specific functions by calling an application’s internal services without UI interaction. This method can alter application state or retrieve data directly, as shown in Fig. 2(a).
@@
The agent-generated shortcuts are created dynamically by the agent itself. They are typically formed by identifying and abstracting repetitive subtasks from the agent’s execution history into new, executable routines. This capability allows the agent to learn from experience and progressively optimize its performance by creating customized shortcuts.
Figure 2: Functional Comparison of APIs, Deep-Links, and RPA Scripts. The figure uses the Amazon app as an example: (a) the open_cart() API call directly opens the shopping cart; (b) the search_product() deep-link directly performs a product search; and (c) an RPA script combines APIs, deep links, and GUI operations to automate a complete workflow.
@@ 3.3 Shortcut Knowledge Base
To validate the feasibility of hybrid GUI-shortcut operations on mobile devices, we construct a predefined knowledge base of commonly used mobile shortcuts. Based on 11 popular apps, it includes 11 APIs, 70 Deep Links, and 7 custom RPA scripts, for a total of 88 predefined shortcuts. The knowledge base provides agents with a rich, queryable resource to select appropriate actions based on task instructions. To ensure real-world relevance, APIs were sourced from official documentation and static analysis of application packages. At the same time, Deep Links were identified by analyzing each application’s declared URI schemes and entry points. The custom RPA scripts were designed to encapsulate common, high-repetition sub-tasks, testing an agent’s ability to leverage complex, pre-built automation routines. More details on the collection and design methodology are available in the Appendix A.3.
@@
Besides predefined shortcuts, MAS-Bench supports evaluating agent-generated shortcuts to test an agent’s learning and abstraction capabilities. In order to demonstrate the effectiveness of MAS-Bench in evaluating agents’ capabilities of generating shortcuts, we have incorporated several contrasting shortcut generation methods for evaluation. We include Macro-level action trajectory replay (abbreviated as action replay) shortcuts that record and replay entire task or sub-task execution trajectories. We also incorporate dynamic shortcuts that require the agent to perform real-time grounding based on execution history, as well as shortcuts generated by the MobileAgent-E framework. This diverse set of generation strategies allows for a comprehensive assessment of an agent’s ability to autonomously create efficient workflows.
@@ 3.4 Tasks Design for GUI-Shortcut Hybrid Agent
MAS-Bench comprises 139 complex tasks derived from real-world scenarios across 11 Android applications, each featuring one or more invocable predefined shortcuts. These tasks reflect authentic user needs, with an average of 9.27 steps for single-app and 17.66 for cross-app workflows based on human operation. A detailed list of the tasks is provided in Appendix A.4.
@@
Figure 3: Evaluation Workflow for Agents’ Shortcut Generation Capability. The process consists of two stages: (a) Shortcut Generation Stage, where the agent creates its shortcut knowledge base; and (b) Quality Evaluation Stage. The Agent-generated shortcuts is imported into a baseline agent for performance testing. The quality of the Agent-generated shortcuts is then measured by comparing this performance against the GUI-only baseline agent and the baseline agent with predefined shortcuts.
To test the agent’s learning and generation capabilities, we introduce scenarios with repetitive sub-tasks. In these scenarios, the knowledge base intentionally lacks predefined shortcuts for these recurrent operations. This design aims to test whether an agent, after completing a sub-task via GUI operations for the first time, can identify the repetitive pattern and then autonomously generate an efficient shortcut for subsequent use.

### v2

@@ Abstract
Shortcuts such as APIs and deep-links have emerged as efficient complements to flexible GUI operations, fostering a promising hybrid paradigm for MLLM-based mobile automation. However, systematic evaluation of GUI–shortcut hybrid agents remains largely underexplored. To bridge this gap, we introduce MAS-Bench, a benchmark that pioneers the evaluation of GUI-shortcut hybrid agents with a specific focus on the mobile domain. Beyond merely using predefined shortcuts, MAS-Bench assesses an agent’s capability to autonomously generate shortcuts by discovering and creating reusable, low-cost workflows. It features 139 complex tasks across 11 real-world applications, a knowledge base of 88 predefined shortcuts (APIs, deep-links, RPA scripts), and 9 evaluation metrics. Experiments demonstrate that hybrid agents achieve up to 68.3% success rate and 39% greater execution efficiency than GUI-only counterparts. Furthermore, our evaluation framework effectively reveals the quality gap between predefined and agent-generated shortcuts, validating its capability to assess shortcut generation methods. MAS-Bench addresses the lack of systematic benchmarks for GUI-shortcut hybrid mobile agents, providing a foundational platform for future advancements in creating more efficient and robust intelligent agents. Project page: https://pengxiang-zhao.github.io/MAS-Bench.
MAS-Bench: A Unified Benchmark for Shortcut-Augmented
Hybrid Mobile GUI Agents
Pengxiang Zhao1thanks: Equal contribution., Guangyi Liu [^1] footnotemark: 1, YaoZhen Liang [^1] footnotemark: 1, Weiqing He1, Zhengxi Lu1, WenHao Wang1, Yuehao Huang1, Yuxiang Chai2, Zhaolu Kang3, Yaxuan Guo2, Hao Wang2, Kexin Zhang1thanks: Corresponding author., Liang Liu2thanks: Project Lead., Yong Liu122footnotemark: 2 1Zhejiang University 2vivo AI Lab 3Peking University
@@ 1 Introduction
The rise of Large Language Models (LLMs) OpenAI (2024); Huang et al. (2025b); Jiang et al. (2025a); Li et al. (2026) is driving the development of Graphical User Interface (GUI) Agents, enabling them to operate diverse digital platforms such as computers, web browsers, and smartphones Xu et al. (2024); Liu et al. (2025b, a). Early research on mobile agents primarily focused on replicating human-like flexibility through GUI-only interaction Wang et al. (2025a); Lu et al. (2026); Xiao et al. (2026); Lu et al. (2025). This approach grants agents the generality to operate on any application, but it often overlooks the significant efficiency advantages offered by more direct methods.
As GUI agents are increasingly applied to complex tasks, enhancing their efficiency has become a central research focus. In this context, hybrid paradigms have emerged as a promising solution, demonstrating effectiveness across a wide range of platforms Zhang et al. (2025a); Guo et al. (2026). This approach combines the speed and reliability of “shortcuts” such as Application Programming Interface (API) calls, deep links, and Robotic Process Automation (RPA) scripts Agostinelli et al. (2022), with the flexibility of GUI operations. As Fig. 1 illustrates, a hybrid agent bypasses multi-step GUI operations by invoking a single shortcut, drastically reducing operational complexity and time.
@@
However, despite these advancements, a framework for systematically evaluating and benchmarking hybrid agents is still lacking Zhang et al. (2025a); Wang et al. (2025b); Huang et al. (2025a). This gap leaves the full potential of GUI-Shortcut agents far from being thoroughly explored and assessed. Mobile platforms provide a controlled and reproducible setting, making them a natural starting point for GUI-shortcut evaluation.
Therefore, we present MAS-Bench, the first benchmark explicitly designed to evaluate GUI-shortcut hybrid mobile agents. The tasks within MAS-Bench are solvable entirely through GUI interactions, but shortcuts can substantially improve execution efficiency. MAS-Bench evaluates whether agents can reliably choose between GUI actions and shortcuts under dynamic mobile environments. Our experiments show that MAS-GLM-4.5V achieves the best success rate of 68.3%. In a matched comparison, MAS-MobileAgent improves over its GUI-only counterpart from 35.2% to 63.3% SR, yielding a 79.8% relative gain and reducing the successful-task step ratio by 38.9%.
Furthermore, we introduce a novel framework to evaluate agents’ capacity to generate new shortcuts from interaction. This framework integrates agent-generated shortcuts into a standard baseline agent and measures subsequent task performance. Our findings reveal a performance gap: while our predefined shortcuts prove highly reliable (100% success rate), agent-generated shortcuts lag in robustness and efficiency. Dynamic shortcuts demonstrate significant potential. Despite their task completion rate of 38% (baseline 43%), they offer the highest efficiency, highlighting them as a promising direction for future research into robust shortcut generation.
Our contributions are threefold:
  • We introduce MAS-Bench, the first benchmark for systematically evaluating GUI-shortcut hybrid mobile agents. It comprises 139 complex tasks spanning 11 real-world applications, supported by a knowledge base of 88 predefined shortcuts and 9 distinct evaluation metrics.
  • We establish extensive baselines across agentic workflows, general-purpose models, and specialized GUI models, demonstrating that GUI-shortcut hybrid operation substantially improves success rate and efficiency while exposing shortcut misuse.
  • We propose the first framework to evaluate an agent’s ability to generate shortcuts from interaction trajectories. Using predefined structural shortcuts as a reference upper-bound baseline, our experiments show that current generated shortcuts still lag behind predefined ones in efficiency and robustness, highlighting a key direction for future work.
@@ 2 Related Work
Mobile task automation evolves from methods based on predefined scripts to more intelligent and adaptive agents driven by LLMs Li et al. (2025); Wang et al. (2025c). Traditional approaches, such as API calls, deep links, and Robotic Process Automation (RPA) scripts, offer direct execution paths but suffer from significant limitations, including invalidation from app updates, and a sensitivity to UI changes that hinders their ability to adapt Kennedy and Everett (2011); Xiao et al. (2025).
The rapid development of LLM-based GUI agents significantly overcomes traditional automation challenges by simulating human interaction through visual or multimodal inputs Cheng et al. (2024). However, complete reliance on LLMs for fine-grained GUI interaction presents new problems: step-by-step action for routine tasks leads to inefficiency and increased costs; in complex operations, LLM hallucinations or misinterpretations can cause cumulative errors, affecting task success rates and reliability Wen et al. (2024). Zhang et al. Zhang et al. (2025a) provide a detailed analysis of API and GUI agents; the former gain attention for their stability and programming integration capabilities, while the latter possesses strong adaptability. These considerations regarding the efficiency and robustness of GUI agents lead to an exploration of how to combine LLM with efficient shortcut methods, thereby promoting the rise of research in GUI-Shortcut agents.
@@
Researchers explore GUI-Shortcut methods to enhance agent capabilities and address the persistent challenges in task completion rate, costs, and operational efficiency for LLM-based GUI agents. For instance, UFO2 Zhang et al. (2025b), a desktop AgentOS, demonstrates robust task execution through its multi-agent architecture and a unified GUI-API action layer that deeply integrates with operating system and application features. Inspired by such work, AppAgentX Jiang et al. (2025b) improves efficiency by evolving high-level actions into shortcuts based on task execution history. MobileAgent-E Wang et al. (2025d) uses a self-evolving module to learn and store general-purpose Tips and reusable Shortcuts from past experiences, continuously improving performance and efficiency over time. While these efforts focused on the mobile domain are promising, they also highlight the urgent need for a systematic method to evaluate the effectiveness of these emerging hybrid operations on mobile devices.
Mobile GUI Agent Benchmark.
As agents evolve from GUI-based interaction to hybrid GUI-Shortcut operations, they place new demands on benchmarking. Among existing benchmarks, MobileAgentBench Wang et al. (2024b) focuses on the foundational GUI navigation and task completion capabilities of LLM-powered mobile agents. AndroidWorld and SPA-Bench Rawles et al. (2024); Chen et al. (2024) benchmark mobile agents on diverse smartphone tasks, spanning dynamic environments, task difficulty. While these benchmarks advance research in various dimensions, there remains a need for a benchmark that can comprehensively evaluate the ability of agents to intelligently discover, decide upon, and execute diverse shortcuts within GUI interactions and also measure the overall task effectiveness of this GUI-Shortcut hybrid operation Zhang et al. (2025b).
@@
AndroidArena Xing et al. (2024) 221
AndroidWorld Rawles et al. (2024) 116
LlamaTouch Zhang et al. (2024) 495
AndroidLab Xu et al. (2024) 138
SPA-Bench Chen et al. (2024) 340
@@ 3.1 Online Evaluation Environment
MAS-Bench is built on a dynamic Android platform to evaluate mobile agent performance on complex tasks comprehensively. Unlike static benchmarks, which are often limited to fixed datasets, our platform enables the real-time assessment of agents’ capacity to select and utilize shortcuts intelligently. To guarantee reproducibility, we employ a snapshot-based reset mechanism that restores the environment to a consistent initial state after each task, and utilize dedicated test accounts to isolate real-world side effects (detailed in Appendix A.2).
@@ 3.2 GUI-Shortcut Hybrid Action Space
To complete tasks, agents in MAS-Bench operate within a hybrid action space that combines conventional GUI interactions with diverse programmatic shortcuts. This space comprises two action types:
@@
  • API: An intent-based programmatic shortcut that invokes application functionality without GUI interaction, as shown in Fig. 2(a).
@@
The agent-generated shortcuts are created dynamically by the agent itself. They are typically formed by identifying and abstracting repetitive subtasks from the agent’s execution history into new, executable routines. This capability allows the agent to learn from experience and progressively optimize its performance across future tasks by creating customized shortcuts.
Figure 2: Functional Comparison of APIs, Deep-Links, and RPA Scripts. The figure uses the Amazon app as an example: (a) the open_cart() API directly opens the shopping cart; (b) the search_product() deep-link directly performs a product search; and (c) an RPA script combines APIs, deep links, and GUI operations to automate a complete workflow.
@@ 3.3 Shortcut Knowledge Base
To validate the feasibility of hybrid GUI-shortcut operations on mobile devices, we construct a predefined knowledge base of commonly used mobile shortcuts. Based on 11 popular apps, it includes 11 APIs, 70 Deep Links, and 7 custom RPA scripts, for a total of 88 predefined shortcuts. The distribution of shortcuts across applications varies based on app complexity and functionality, ranging from 2 shortcuts for Google Calendar to 21 for Fitbit (see Appendix Table 17). To ensure real-world relevance, APIs were identified from official documentation and static analysis of application packages. These APIs target intent-accessible endpoints and are executed through standard Android IPC/ADB commands without requiring root access, debug signatures, or emulator system modifications. At the same time, Deep Links were identified by analyzing each application’s declared URI schemes and entry points. The custom RPA scripts were designed to encapsulate common, high-repetition sub-tasks, testing an agent’s ability to leverage complex, pre-built automation routines. More details on the collection and design methodology are available in the Appendix A.4.
@@
Besides predefined shortcuts, MAS-Bench supports evaluating agent-generated shortcuts to test an agent’s learning and abstraction capabilities. In order to demonstrate the effectiveness of MAS-Bench in evaluating agents’ capabilities of generating shortcuts, we have incorporated several contrasting shortcut generation methods for evaluation. We include Macro-level action trajectory replay (abbreviated as action replay) shortcuts that record and replay entire task or sub-task execution trajectories. We also incorporate dynamic shortcuts that require the agent to perform real-time grounding based on execution history, as well as shortcuts generated by the MobileAgent-E Wang et al. (2025d). This diverse set of generation strategies allows for a comprehensive assessment of an agent’s ability to autonomously create efficient workflows.
@@ 3.4 Tasks Design for GUI-Shortcut Hybrid Agent
MAS-Bench comprises 139 complex tasks derived from real-world scenarios across 11 Android applications, each featuring one or more invocable predefined shortcuts. These tasks reflect authentic user needs, with an average of 9.27 steps for single-app and 17.66 for cross-app workflows based on human operation. Tasks are categorized into three difficulty levels (detailed in Appendix A.5), with category-level examples summarized in Table 6.
@@
Figure 4: Evaluation Workflow for Agents’ Shortcut Generation Capability. The process consists of two stages: (a) Shortcut Generation Stage, where the agent creates its shortcut knowledge base; and (b) Quality Evaluation Stage. The agent-generated shortcut knowledge base is imported into a baseline agent for performance testing. Its quality is then measured by comparing this performance against the GUI-only baseline agent and the baseline agent with predefined shortcuts.
To test the agent’s learning and generation capabilities, we introduce scenarios with repetitive sub-tasks. In these scenarios, the knowledge base intentionally lacks predefined shortcuts for these recurrent operations. This design aims to test whether an agent, after completing a sub-task via GUI operations for the first time, can identify the repetitive pattern and then autonomously generate an efficient shortcut for subsequent use. Notably, some tasks in MAS-Bench are designed as incremental variations, where a task builds upon the previous one by adding additional requirements. This incremental design serves two purposes: it creates realistic scenarios where agents must handle evolving user needs, and it provides an ideal testbed for evaluating shortcut generation by presenting repeated sub-task patterns that agents should learn to abstract into reusable shortcuts.